OpenStudy (anonymous):

Factor \[ a^{7}+a^{2}+1 \] and deduce that 1280000401 is composite.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\rightarrow a^7 + a^2 + 1^2 }\) Try to factor it.

OpenStudy (lgbasallote):

did you get that @eliassaab ?

Parth (parthkohli):

Think about how yo may factor it.

OpenStudy (anonymous):

I know how to do it. I posted it just for fun.

OpenStudy (callisto):

eliassaab is a retired maths professor... obviously... this question is for fun only :|

OpenStudy (anonymous):

Hint: Evaluate this quotient \[\frac{a^7+a^2+1}{a^2+a+1} \] using long division.

OpenStudy (cwrw238):

i did the long division on paper ( awkward on computer) a^2 + a + 1 is a factor the quotient is a^5 - a^4 + a^2 - a + 1

OpenStudy (cwrw238):

not sure if that will factor

OpenStudy (callisto):

(1+a+a^2) (1-a+a^2-a^4+a^5) <- That's it!~

OpenStudy (cwrw238):

i'm not sure how to do the second part however i made a guess that 'a' would a number ending in one (pure intuition) i tried ten but the 10^7 + 10^2 + 1 fell short of the quoted number tried 20 and it worked out 20^2 + 20 + 1 = 421 this divided into 1280000401 exactly giving 3040381 which = 20^5 - 20^4 + 20^2 - 20 + 1

OpenStudy (cwrw238):

* correction i guessed that 'a' would end in 0.

OpenStudy (anonymous):

I think you got it \[a^7+a^2+1=\left(a^2+a+1\right) \left(a^5-a^4+a^2-a+1\right ) \]

OpenStudy (anonymous):

Now, you can answer the second part.

OpenStudy (anonymous):

\[20^7+20^2+1=1280000401\]